OpenStudy (anonymous):

the inverse function of (3x-8)/(7x+9) give the domain and range as well... @Calculus1

OpenStudy (anonymous):

\[\frac{3x - 8}{7x + 9}\] The domain of this is \[(-\infty, \frac{9}{7} \cup (\frac{9}{7}, \infty)\] And range is \[(-\infty, \infty)\] This will make finding the domain and range of the inverse much eaiser. now to find the inverse you first write it like this \[y = \frac{3x - 8}{7x + 9}\] Now switch the x's and y'x \[x = \frac{3y - 8}{7y + 9}\] solve for y

OpenStudy (anonymous):

The domain of the inverse function is the range of the original function. And the range of the inverse function is the domain of the regular function.

OpenStudy (anonymous):

\[(-\infty, - \frac{9}{7}) \cup (-\frac{9}{7},\infty)\] This is should be the domain of the regular function.